# Error-Correcting Codes for Two Bursts of $t_{1}$-Deletion-$t_{2}$-Insertion with Low Computational Complexity

### v1

@@ V-B Two Bursts of $(t_{1},t_{2})$-DI ECCs
Case 2. Lemma 7-b) holds. In this case, we can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$. Following (35), we know that there exist an interval $\mathcal{L}_{j}$ such that $\mathcal{J}\subseteq\mathcal{L}_{j}$, where $j\in[\lceil n/\rho_{2}\rceil-1]$. Then, the proof is similar to that of subcase 2) in Case 1, with the only difference being the use of $g^{(i)}(\mathbf{x})$ in place of $h^{(i)}(\mathbf{x})$. The details are therefore omitted.
Following Lemma 3, the equations (36)-(38), the length of $\psi(\mathbf{x})$ satisfies
@@ V-C Complexity
In Construction 1, two bursts of $(t_{1},t_{2})$-DI errors in a binary sequence $\mathbf{x}$ are corrected by the syndrome compression technique directly. Regarding to Construction 2, we firstly locate two bursts of $(1,t^{\prime}-1)$-DS in the first row of matrix representation of $\mathbf{x}$ using a function achieved by the syndrome compression technique, and then correct two bursts of $(t_{1},t_{2})$-DI of $\mathbf{x}$. As shown in [24], the main computational cost of the syndrome compression technique arises from a brute-force search for a parameter
@@
Moreover, for the case of $t_{1}>t_{2}$, the functions $\phi^{(a)}_{i}(\mathbf{x})$, $h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ in Construction 2 are computable in time $O((2\rho_{1})^{4})$, $O((2\rho_{1})^{2})$ and $O((2\rho_{2})^{4})$, where $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$, and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$, respectively. This means that the functions $\phi^{(a)}_{i}(\mathbf{x}),h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ have at most $O({(\log n)}^{4})$ time complexity. Hence, the computational complexity of Construction 2 is mainly from $f(X[1])$, i.e., the cardinality of $\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}(\mathbf{x})$. Consequently, comparing (54) to (55) and (56), the Construction 2 reduces computational complexity by a factor of $(t_{1}-t_{2})^{4}$.
In addition, for the case of $t_{1}=t_{2}$, two bursts of $(t_{1},t_{2})$-DI errors are transferred to four substitutions, which are corrected by a GRS code. Both of the encoding and decoding complexities for GRS codes are $O(n\log n)$, which is extremely less than $O(n^{4})$.
@@
$|\mathcal{N}_{2,(10,2)}(\mathbf{x})|={238\choose 2}{241\choose 2}\times 2^{16}\approx 2^{46},$ (57)
@@ 12 $256$ $2^{46}$ $2^{18}$
13 $\left|\mathcal{N}_{2,(t_{1},t_{2})}(\mathbf{x})\right|$ Operations of GRS codes
@@ 16 $512$ $(2,2)$ $2^{34}$ $2^{19}$
Furthermore, a numerical comparison between the codes in Theorems 5 and 8 for several fixed parameters is presented in Table I. It can be observed that, across all parameter settings, the computational complexity of those in Theorem 8 is extremely lower than the one in Theorem 5. For instance, when $n=1024,512,256$, if there are two bursts of $(t_{1}=10,t_{2}=6)$-DI occur in a binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, the codes in Theorem 8 yield a significant reduction in the size of $\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})$, improving from $\left|\mathcal{N}^{DI}_{2,(10,6)}(\mathbf{x})\right|\approx 2^{62},2^{58}$ and $2^{54}$ in Theorem 5 to $\left|\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})\right|\approx 2^{34},2^{30}$ and $2^{26}$, respectively, a constant factor improvement of $2^{28}$ across all code lengths (Rows 7-9).

### v2

@@ V-B Two Bursts of $(t_{1},t_{2})$-DI ECCs
Case 2. Lemma 7-b) holds. In this case, we can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime}+1)(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$. Following (35), we know that there exist an interval $\mathcal{L}_{j}$ such that $\mathcal{J}\subseteq\mathcal{L}_{j}$, where $j\in[\lceil n/\rho_{2}\rceil-1]$. Then, the proof is similar to that of subcase 2) in Case 1, with the only difference being the use of $g^{(i)}(\mathbf{x})$ in place of $h^{(i)}(\mathbf{x})$. The details are therefore omitted.
Following Lemma 3, and Equations (36)-(38), the length of $\psi(\mathbf{x})$ satisfies
@@ V-C Complexity
In Construction 1, two bursts of $(t_{1},t_{2})$-DI errors in a binary sequence $\mathbf{x}$ are corrected by a direct application of the syndrome compression technique. Regarding to Construction 2, we firstly locate two bursts of $(1,t^{\prime}-1)$-DS in the first row of matrix representation of $\mathbf{x}$ using a function achieved by the syndrome compression technique, and then correct two bursts of $(t_{1},t_{2})$-DI of $\mathbf{x}$. As shown in [23], the main computational cost of the syndrome compression technique arises from a brute-force search for a parameter
@@
Moreover, for the case of $t_{1}>t_{2}$, the functions $\phi^{(a)}_{i}(\mathbf{x})$, $h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ in Construction 2 are computable in time $O((2\rho_{1})^{4})$, $O((2\rho_{1})^{2})$ and $O((2\rho_{2})^{4})$, where $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime}+1)(t_{1}-t_{2})$, and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime}+1)(t_{1}-t_{2})$, respectively. This means that the functions $\phi^{(a)}_{i}(\mathbf{x}),h^{(i)}(\mathbf{x})$ and $g^{(i)}(\mathbf{x})$ have at most $O({(\log n)}^{4})$ time complexity. Hence, the computational complexity of Construction 2 is mainly from $f(X[1])$, i.e., the cardinality of $\mathcal{N}^{DS}_{2,(1,t^{\prime}-1)}(\mathbf{x})$. Consequently, comparing (54) to (55) and (56), the codes in Construction 2 reduces computational complexity by a factor of $(t_{1}-t_{2})^{4}$.
In addition, for the case of $t_{1}=t_{2}$, two bursts of $(t_{1},t_{2})$-DI errors are transferred to four substitutions, which are corrected by a coset of a RS code. Both of the encoding and decoding complexities for RS codes are $O(n\log n)$, which is extremely less than $O(n^{4})$.
@@
$|\mathcal{N}_{2,(10,2)}^{DI}(\mathbf{x})|={238\choose 2}{241\choose 2}\times 2^{16}\approx 2^{46},$ (57)
@@ 12 $256$ $2^{46}$ $2^{18}$
13 Operations of RS codes
@@ 16 $512$ $(2,2)$ $2^{34}$ $2^{19}$
Furthermore, a numerical comparison between the codes in Theorems 5 and 8 for several fixed parameters is presented in Table I. It can be observed that, across all parameter settings, the computational complexity of those in Theorem 8 is extremely lower than the one in Theorem 5. For instance, when $n=1024,512,256$, if there are two bursts of $(t_{1}=10,t_{2}=6)$-DI occur in a binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$, the codes in Theorem 8 yield a significant reduction in the size of $\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})$, improving to $\left|\mathcal{N}^{DS}_{2,(1,2)}(\mathbf{x})\right|\approx 2^{34},2^{30}$ and $2^{26}$ from $\left|\mathcal{N}^{DI}_{2,(10,6)}(\mathbf{x})\right|\approx 2^{62},2^{58}$ and $2^{54}$ in Theorem 5, respectively, a constant factor improvement of $2^{28}$ across all code lengths (Rows 7-9).